# Activation Approximations Can Incur Safety Vulnerabilities Even in Aligned LLMs: Comprehensive Analysis and Defense

### v1

@@ 3.3 Activation Approximation Error
$\texttt{MLP}_{l}(\theta_{l}|\mathbf{e}_{l})=(\alpha(\mathbf{e}_{l}\cdot\mathbf{W}^{\text{up}}_{l}))\cdot\mathbf{W}^{\text{down}}_{l},$
@@
Table 2: Benchmark of different training-free activation approximation techniques on Llama-3.1-8B-Instruct. Note that the baseline PPL is 8.05, ASR is 0.19.
[…]
Activation Polynomialization Iron[19] MPC $\mathcal{N}(0,0.064^{2})$ Lap(0, 0.049) 43.74 58.07
BOLT [21] MPC $\mathcal{N}(0,0.042^{2})$ Lap(0, 0.036) 15.81 35.76
BumbleBee [22] MPC $\mathcal{N}(0,0.026^{2})$ Lap(0, 0.018) 11.28 20.56
NEXUS [23] FHE $\mathcal{N}(0,0.031^{2})$ Lap(0, 0.014) 12.93 21.15
Activation Sparsification TEAL [24] 10% sparsity $\mathcal{N}(0,0.35^{2},|X|\leq 0.04)$ $\text{Lap}(0,0.024,|X|\leq 0.003)$ 8.05 5.00
25% sparsity $\mathcal{N}(0,0.35^{2},|X|\leq 0.11)$ $\text{Lap}(0,0.024,|X|\leq 0.007)$ 8.09 28.85
50% sparsity $\mathcal{N}(0,0.35^{2},|X|\leq 0.24)$ $\text{Lap}(0,0.024,|X|\leq 0.017)$ 8.57 57.12
90% sparsity $\mathcal{N}(0,0.35^{2},|X|\leq 0.57)$ $\text{Lap}(0,0.024,|X|\leq 0.055)$ 206.47 0.02
Activation Quantization SmoothQuant[27] W16A8 $\mathcal{N}(0,0.027^{2})$ Lap(0,0.019) 9.56 41.92
W16A4 $\mathcal{N}(0,0.035^{2})$ Lap(0,0.024) 10.35 55.19
OmniQuant[28] W16A8 $\mathcal{N}(0,0.029^{2})$ Lap(0,0.028) 9.85 33.46
W16A4 $\mathcal{N}(0,0.036^{2})$ Lap(0,0.037) 10.79 47.11
@@
$\epsilon^{up}=\mathcal{F}^{\text{LayerNorm}}(x)-\mathcal{F}^{\text{LayerNorm}}_{\text{approx}}(x).$
@@
Approximation Error of Activation Sparsification Following the definition of activation sparsification in TEAL[24], for a vector $\mathbf{X}=(x_{1},...,x_{n})$ and sparsity level $p\in[0,1]$, define the threshold $t$ as:
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{P}(|x_{i}|\leq t)=p$
The sparsification function $S_{t}(\cdot):\mathbb{R}\rightarrow\mathbb{R}$ is defined as:
@@
$=\text{round}(c^{\text{FP16}}\cdot x^{\text{FP16}})=x^{\text{INT4}}.$
where $c^{\text{FP16}}$ is the quantization scale. Dequantization is the inverse:
@@
Distribution Fitting for Approximation Errors We summarize approximation errors from different activation approximations by introducing two noise components $\epsilon^{up}_{l},\epsilon^{down}_{l}$, which is formalized by the following definition.
@@
$\texttt{MLP}_{l}^{\epsilon}(\theta_{l}|\mathbf{e}_{l})=(\alpha((\mathbf{e}_{l}+\epsilon^{up}_{l})\cdot\mathbf{W}^{\text{up}}_{l})+\epsilon^{down}_{l})\cdot\mathbf{W}^{\text{down}}_{l},$
@@
We then fit the error distributions of $\epsilon^{up}_{l},\epsilon^{down}_{l}$ for representative methods from all three activation approximation techniques. The errors are recorded from activations collected using Llama-3.1-8B-Instruct on the Alpaca dataset [96]. Figure 2 presents the resulting error distributions, where $\epsilon^{up}_{l}$ and $\epsilon^{down}_{l}$ are zero-mean and unimodal, falling into two distinctly shaped distributions. The activation error $\epsilon^{up}_{l}$ before the $\mathbf{W}^{\text{up}}$ tends to be Gaussian-like distribution, while the activation error $\epsilon^{down}_{l}$ before the $\mathbf{W}^{\text{down}}$ tends to be Laplacian-like distribution.
Figure 3: Approximation of activation before $\mathbf{W}^{\text{up}}$ on different safety-aligned LLMs.
Figure 4: Approximation of activation before $\mathbf{W}^{\text{down}}$ on different safety-aligned LLMs.
Family Model Sensitive Baseline Error Before $\mathbf{W}^{\text{up}}$ Error Before $\mathbf{W}^{\text{down}}$
Layers ASR PPL mva ASR PPL mva ASR PPL
Llama Llama-2-7B-Chat [0-3] 1.34 7.84 $\mathcal{N}(0,0.045^{2})$ 56.73 7.91 Lap(0, 0.100) 49.80 10.96
Llama-2-13B-Chat [0-3] 0.00 7.07 $\mathcal{N}(0,0.042^{2})$ 47.88 8.05 Lap(0, 0.125) 52.12 18.46
Llama-3.1-8B-Instruct [0-3] 0.19 8.05 $\mathcal{N}(0,0.075^{2})$ 69.23 14.79 Lap(0, 0.085) 67.50 11.70
Phi3 Phi-3-Mini-4K-Instruct [0-3] 0.00 6.71 $\mathcal{N}(0,0.040^{2})$ 82.61 24.79 Lap(0, 0.120) 32.38 9.46
Phi-3.5-Mini-Instruct [0-3] 0.77 6.86 $\mathcal{N}(0,0.033^{2})$ 79.50 21.43 Lap(0, 0.100) 44.72 7.45
Mistral Mistral-7B-Instruct-v0.3 [0-6] 34.61 6.13 $\mathcal{N}(0,0.200^{2})$ 73.18 6.40 Lap(0, 0.075) 71.02 6.24
Mixtral-8x7B-Instruct-v0.1 [0-4] 1.15 4.61 $\mathcal{N}(0,0.400^{2})$ 82.47 12.09 Lap(0, 0.225) 87.64 19.31
Zephyr-7B-$\beta$ [0-4] 22.31 7.01 $\mathcal{N}(0,0.250^{2})$ 88.30 7.76 Lap(0, 0.113) 82.35 9.07
Qwen Qwen2-7B-Instruct [0-4] 0.38 8.46 $\mathcal{N}(0,0.300^{2})$ 52.16 14.01 Lap(0, 0.058) 73.82 23.94
Qwen2.5-32B-Instruct [0-4] 0.00 5.74 $\mathcal{N}(0,0.200^{2})$ 37.48 5.81 Lap(0, 0.043) 46.13 21.54
@@ 3.4 Further Analysis and Key Observations
Since errors were introduced in the outputs of activation functions as shown in the case studies, we conduct continuous noise simulation experiments on 10 safety-aligned LLMs to explore the impact of activation approximations on model safety and utility. Based on the noise distributions observed in Figure 2, we apply different levels of approximation to the activation functions of commonly used safety-aligned LLMs. The results are shown in Figure 3 and Figure 4.
We collect the corresponding ASRand Perplexity scores produced by the models after noises have been added to their activation outputs before $\mathbf{W}^{\text{up}}$ and $\mathbf{W}^{\text{down}}$. The $x$-axis is the standard deviations of the noise, the left $y$-axis represents the ASRand the right $y$-axis measures the PPL scores, respectively, in different scales.
We can see that there exists a general trend where as the simulated approximation errors increase, ASR gradually grows at an increasing rate until it reaches a peak at a certain noise level. Throughout this period, the PPL scores appear to not change much and stay at a relatively low value. This means that during this interval, safety-aligned LLMs will fail to reject the malicious question and produce meaningful replies to malicious requests.
When the noise grows too much, the PPL starts to increase exponentially and the ASR begins to decrease dramatically, at which point the LLM will output meaningless responses like “####”, “I’m I’m I’m”. At this point, the model becomes unusable both to the normal users and to the attackers.
Since LLMs are observed to lose safety before they lose their utility as the noise grows, these models expose a jailbreak vulnerability within a certain approximation interval. Consequently, we can find the most vulnerable approximation (denoted as mva) that maximizes the ASR of the model.
@@
where $\texttt{HarmCLS}(\cdot)$ is a harmful behaviour classifier, $\mathcal{D}$ is the harmful prompt dataset, and we use Advbench [46] in our experiments, $\pi^{\epsilon}_{\theta}(\cdot|x)$ refers to the response at inference $x$ given by LLM with parameter $\theta$ and activation approximation noise with magnitude $\epsilon$.
We record the mva of each model in Table 3 as well as the ASR and PPL under this approximation, which will serve as a guideline for designing the defense method in Section 4.
@@
Activation approximation errors in the first few layers are the most detrimental to safety, while the approximations in the later layers have minimal impact on safety.
Assuming that $\theta$ is an $L$-layer model $(\theta_{1},\theta_{2},...\theta_{L})$, for and input $x$, the inference with activation approximation can be defined as
$\pi^{\epsilon}_{\theta}(\cdot|x)$ $=\bm{f}_{L}(\texttt{MLP}_{L}^{\epsilon_{L}}(\theta_{L}|\mathbf{e}_{L}))\circ\dots\circ\bm{f}_{1}(\texttt{MLP}_{1}^{\epsilon_{1}}(\theta_{1}|\mathbf{e}_{1})),$
$\mathbf{e}_{1}$ $=\texttt{Attention}(\mathcal{T}(x)),$
where $\circ$ represents the layer-wise concatenation of LLM, and $\mathcal{T}(x)$ is the tokenizer function, $\mathbf{e}_{1}$ is the first-layer embedding input to MLP. We provide an analysis of the impact of layer-wise approximations on the safety of LLMs, to explore which layers lack robustness to approximations. Define the inference for layer-wise approximations as:
$\pi^{\epsilon}_{\theta}(\cdot|x)_{k}$ $=\underbrace{\bm{f}_{L}(\texttt{MLP}_{L}(\theta_{L}|\mathbf{e}_{L}))\circ\dots\circ\bm{f}_{k+1}(\texttt{MLP}_{k+1}(\theta_{k+1}|\mathbf{e}_{k+1}))}_{\text{The last $L-k$ layers are not approximate}}$
[…]
Figure 6 shows the ASR of LLMs after adding mva noise on different layers. Note that $\pi^{\epsilon}_{\theta}(\cdot|x)_{k}$ indicates that the approximation noise is added after the $k$-th layer, e.g., $\pi^{\epsilon}_{\theta}(\cdot|x)_{1}$ indicates that the activations of all layers are approximated, and $\pi^{\epsilon}_{\theta}(\cdot|x)_{L}$ indicates that only the activations in the last layer is approximated.
Based on Figure 6, we observe that only the first few layers are the most fatal, indicating that the first few layers lack robustness to activation approximations, which we call “sensitive layers”. We also record the sensitive layers of each model in Table 3, which are the layers we focus on when designing the defense strategy.
@@
Figure 6: Layer-wise activation approximations on different safety-aligned LLMs. The green lines indicate approximation noise added to activations before $\mathbf{W}^{\text{up}}$. The orange lines indicate approximation noise added to activations before $\mathbf{W}^{\text{down}}$.
@@
We can observe that before activation approximations, there exists a clear distinction between benign and harmful activations, with an average cosine similarity of $0.52$ among harmful activations. After activation approximations, harmful prompts become more scattered and diverse, overlapping significantly with benign regions, resulting in an average cosine similarity of $0.40$ among harmful activations. This shift caused by activation approximations enables harmful prompts to generate model internal states that resemble those of benign prompts, thereby bypassing safety checks.

### v2

@@ 3.3 Activation Approximation Error
Table 3: Benchmark of different training-free activation approximation techniques on Llama-3.1-8B-Instruct. Note that the baseline ASR is 0.19, HS is 1.05, PPL is 8.05 and MMLU is 73.2.
[…]
Activation Polynomialization Iron[18] MPC $\mathcal{N}(0,0.064^{2})$ Lap(0, 0.049) 58.07 4.42 43.74 55.4
BOLT [20] MPC $\mathcal{N}(0,0.042^{2})$ Lap(0, 0.036) 35.76 3.57 15.81 62.3
BumbleBee [21] MPC $\mathcal{N}(0,0.026^{2})$ Lap(0, 0.018) 20.56 2.95 11.28 65.0
NEXUS [22] FHE $\mathcal{N}(0,0.031^{2})$ Lap(0, 0.014) 21.15 3.04 12.93 63.9
Activation Sparsification TEAL [23] 10% sparsity $\mathcal{N}(0,0.35^{2},|X|\leq 0.04)$ $\text{Lap}(0,0.024,|X|\leq 0.003)$ 5.00 2.17 8.05 73.0
25% sparsity $\mathcal{N}(0,0.35^{2},|X|\leq 0.11)$ $\text{Lap}(0,0.024,|X|\leq 0.007)$ 28.85 3.29 8.09 72.5
50% sparsity $\mathcal{N}(0,0.35^{2},|X|\leq 0.24)$ $\text{Lap}(0,0.024,|X|\leq 0.017)$ 57.12 4.38 8.57 70.8
90% sparsity $\mathcal{N}(0,0.35^{2},|X|\leq 0.57)$ $\text{Lap}(0,0.024,|X|\leq 0.055)$ 0.02 1.00 206.47 26.5
Activation Quantization SmoothQuant[26] W16A8 $\mathcal{N}(0,0.027^{2})$ Lap(0,0.019) 41.92 3.75 9.56 68.4
W16A4 $\mathcal{N}(0,0.035^{2})$ Lap(0,0.024) 55.19 4.20 10.35 67.5
OmniQuant[27] W16A8 $\mathcal{N}(0,0.029^{2})$ Lap(0,0.028) 33.46 3.41 9.85 68.1
W16A4 $\mathcal{N}(0,0.036^{2})$ Lap(0,0.037) 47.11 4.05 10.79 66.3
@@
$\texttt{MLP}_{l}(\theta_{l}|\mathbf{e}_{l})=(\alpha(\mathbf{e}_{l}\cdot\mathbf{W}^{\text{up}}_{l}))\cdot\mathbf{W}^{\text{down}}_{l},$ (1)
@@
$\epsilon^{up}=\mathcal{F}^{\text{LayerNorm}}(x)-\mathcal{F}^{\text{LayerNorm}}_{\text{approx}}(x);$
@@
(a) Activation polynomialization error, with BOLT [20] as the example.
(b) Activation sparsification error, with TEAL [23] in 50% sparsity.
(c) Activation quantization error, with SmoothQuant [26] in 8-bit.
Figure 2: Activation approximation error distributions on Llama-3.1-8B-Instruct with different techniques. The first column presents the approximation error $\epsilon^{up}$ before multiplication with $\mathbf{W}^{\text{up}}$, and the second column presents the error $\epsilon^{down}$ before multiplication with $\mathbf{W}^{\text{down}}$. Best-fit Gaussian/Laplace distributions are overlaid in black.
Approximation Error of Activation Sparsification Following the definition of activation sparsification in TEAL[23], for a vector $\mathbf{X}=(x_{1},...,x_{n})$ and sparsity level $p\in[0,1]$, define the threshold $t$ as: $\frac{1}{n}\sum_{i=1}^{n}\mathbb{P}(|x_{i}|\leq t)=p.$ The sparsification function $S_{t}(\cdot):\mathbb{R}\rightarrow\mathbb{R}$ is defined as:
@@
$=\text{round}(c^{\text{FP16}}\cdot x^{\text{FP16}})=x^{\text{INT4}},$
where $c^{\text{FP16}}$ is quantization scale. And Dequantization is:
@@
Distribution Fitting for Approximation Errors. We summarize approximation errors from different activation approximations by introducing two noise components $\epsilon^{up}_{l},\epsilon^{down}_{l}$, which is formalized by the following definition.
@@
$\texttt{MLP}_{l}^{\epsilon}(\theta_{l}|\mathbf{e}_{l})=(\alpha((\mathbf{e}_{l}+\epsilon^{up}_{l})\cdot\mathbf{W}^{\text{up}}_{l})+\epsilon^{down}_{l})\cdot\mathbf{W}^{\text{down}}_{l},$ (2)
@@
We then fit the error distributions of $\epsilon^{up}_{l},\epsilon^{down}_{l}$ for representative methods from all three activation approximation techniques. The errors are recorded from activations collected using Llama-3.1-8B-Instruct on the Alpaca dataset [86]. Figure 2 presents the resulting error distributions, where $\epsilon^{up}_{l}$ and $\epsilon^{down}_{l}$ are zero-mean and unimodal, falling into two distinctly shaped distributions. The activation error $\epsilon^{up}_{l}$ before $\mathbf{W}^{\text{up}}$ tends to be a Gaussian distribution, while the activation error $\epsilon^{down}_{l}$ before $\mathbf{W}^{\text{down}}$ tends to be a Laplacian distribution.
Figure 3: Safety and utility performance with different approximation error $\epsilon^{\text{up}}$ (before multiplication with $\mathbf{W}^{\text{up}}$).
Figure 4: Safety and utility performance with different approximation error $\epsilon^{\text{down}}$ (before multiplication with $\mathbf{W}^{\text{down}}$).
@@ 3.4 Further Analysis and Key Observations
Since errors are introduced in the outputs of activation functions as shown in the case studies, we conduct continuous noise simulation experiments on 10 safety-aligned LLMs to explore the impact of activation approximations on model safety and utility. Based on the noise distributions observed in Figure 2, we apply different levels of approximation to the activation functions of commonly used safety-aligned LLMs.
Figure 3 and Figure 4 present results measured by ASRand Perplexity. We can see that there exists a general trend where, as the simulated approximation errors increase, ASR gradually grows at an increasing rate until it reaches a peak at a certain noise level. Throughout this period, the PPL scores appear not to change much and stay at a relatively low value. This means that during this interval, safety-aligned LLMs will fail to reject the malicious question and produce meaningful replies to malicious requests. When the noise grows too much, the PPL starts to increase exponentially and the ASR begins to decrease dramatically, at which point the LLM will output meaningless responses like “####”, “I’m I’m I’m”. At this point, the model becomes unusable both to the normal users and to the attackers. Since LLMs are observed to lose safety before they lose their utility as the noise grows, these models expose a jailbreak vulnerability within a certain approximation interval. Definition 3 below formalizes the activation approximation errors in a layer-wise manner.
Definition 3.
Assuming that $\theta$ is an $L$-layer model $(\theta_{1},\theta_{2},...\theta_{L})$, for the input $x$, the inference with activation approximation errors $\bm{\epsilon}=\{\epsilon_{1},...,\epsilon_{L}\}$ can be defined as
$\pi^{\epsilon}_{\theta}(\cdot|x)=\bm{f}^{\epsilon_{L}}_{L}(\theta_{L}|\mathbf{e}_{L})\circ\bm{f}^{\epsilon_{L-1}}_{L-1}(\theta_{L-1}|\mathbf{e}_{L-1})\dots\circ\bm{f}^{{}^{\epsilon_{1}}}_{1}(\theta_{1}|\mathcal{T}(x)),$ (3)
where $\circ$ represents the layer-wise concatenation of LLM, and $\mathcal{T}(x)$ is the tokenizer function, $\mathbf{e}_{l}$ is the embedding input to $l$-th layer, and $\bm{f}^{\epsilon_{l}}_{l}$ means the $l$-the layer with approximation errors $\epsilon_{l}$ in MLP block, defined in Definition 2.
Model Baseline Error Before $\mathbf{W}^{\text{up}}$ Error Before $\mathbf{W}^{\text{down}}$
ASR HS PPL MMLU mva ASR HS PPL MMLU mva ASR HS PPL MMLU
Llama-2-7B-Chat 1.34 1.14 7.84 45.8 $\mathcal{N}(0,0.045^{2})$ 56.73 3.27 7.91 45.0 Lap(0, 0.100) 49.80 2.99 10.96 43.5
Llama-2-13B-Chat 0.00 1.00 7.07 53.6 $\mathcal{N}(0,0.042^{2})$ 47.88 2.91 8.05 52.1 Lap(0, 0.125) 52.12 3.18 18.46 48.0
Llama-3.1-8B-Instruct 0.19 1.05 8.05 73.2 $\mathcal{N}(0,0.075^{2})$ 69.23 3.78 14.79 66.5 Lap(0, 0.085) 67.50 3.81 11.70 68.0
Phi-3-Mini-4K-Instruct 0.00 1.00 6.71 68.8 $\mathcal{N}(0,0.040^{2})$ 82.61 4.42 24.79 61.9 Lap(0, 0.120) 32.38 2.32 9.46 65.5
Phi-3.5-Mini-Instruct 0.77 1.10 6.86 68.1 $\mathcal{N}(0,0.033^{2})$ 79.50 4.25 21.43 61.3 Lap(0, 0.100) 44.72 2.80 7.45 65.0
Mistral-7B-Instruct-v0.3 34.61 3.15 6.13 61.5 $\mathcal{N}(0,0.200^{2})$ 73.18 3.99 6.40 60.3 Lap(0, 0.075) 71.02 3.97 6.24 60.5
Mixtral-8x7B-Instruct-v0.1 1.15 1.12 4.61 70.6 $\mathcal{N}(0,0.400^{2})$ 82.47 4.41 12.09 63.0 Lap(0, 0.225) 87.64 4.61 19.31 58.3
Zephyr-7B-$\beta$ 22.31 2.90 7.01 61.4 $\mathcal{N}(0,0.250^{2})$ 88.30 4.66 7.76 59.2 Lap(0, 0.113) 82.35 4.42 9.07 58.5
Qwen2-7B-Instruct 0.38 1.06 8.46 70.5 $\mathcal{N}(0,0.300^{2})$ 52.16 3.17 14.01 63.4 Lap(0, 0.058) 73.82 3.96 23.94 60.0
Qwen2.5-32B-Instruct 0.00 1.00 5.74 78.5 $\mathcal{N}(0,0.200^{2})$ 37.48 2.60 5.81 77.2 Lap(0, 0.043) 46.13 2.88 21.54 73.6
We can find the Most Vulnerable Approximation (denoted as mva) that maximizes the ASR of the model.
@@
where $\texttt{HarmCLS}(\cdot)$ is a harmful behaviour classifier, $\mathcal{D}$ is the harmful prompt dataset, and we use Advbench [60] with 520 samples in our experiments.
Table 4 further presents the HS and MMLU metrics, which exhibit trends consistent with the ASR and PPL results. Specifically, under each model’s mva, the utility after activation approximations remains comparable to the baseline, as evidenced by the relatively stable MMLU. However, safety is compromised, as indicated by the increased HS. Notably, the mva values recorded for each model in Table 4 will serve as a reference for the design of the defense method in Section 4.
Figure 5: Safety and utility performance with varying numbers of perturbed layers ($\tau$) in Eq.(6). Note that when $\tau=0$, no activation approximations are applied. When $\tau>0$, $\tau$ layers are perturbed.
@@
Activation approximation errors in the first few layers are the most detrimental to safety, while the approximations in the later layers have minimal impact.
Observation I is mainly based on experimental results. To further study the impact of approximation errors across different layers, we formulate the identification of sensitive layers as an optimization problem: finding $\bm{\epsilon}=\{\epsilon_{1},...,\epsilon_{L}\}$ that maximizes ASR, subject to $l_{0}$-norm constraint to perturb only the sensitive layers. Since Eq.(4) is non-differentiable, it is less suitable for identifying $\bm{\epsilon}$ that maximizes ASR through gradient descent. Following [60], we resort to the negative log probability of harmful target sequences of tokens (i.e., $x^{*}$ represents the phrase “Sure, here is how to build a bomb.”) as the harmful loss:
$\mathcal{L}_{\text{harm}}(\bm{\epsilon})=-\mathbb{E}_{x\sim\mathcal{D}_{\text{BeaverTails}}}\log\pi^{\bm{\epsilon}}_{\theta}(x^{*}|x).$ (5)
We can write the objective in below for the adversarial attack with activation perturbations:
$\operatorname*{minimize}_{\bm{\epsilon}}~{}\mathcal{L}_{\text{harm}}(\bm{\epsilon}),\;\;\text{s.t.}~{}||\bm{\epsilon}||_{0}\leq\tau.$ (6)
The perturbation is $l_{0}$-norm bounded (keep only $\tau$ as non-zero), and the objective is optimized with projected SGD. Under this constraint, small noise will be truncated, leaving non-zero noise corresponding to sensitive layers.
@@
Figure 7: Layer-wise noise distributions $\bm{\epsilon}=\{\epsilon_{1},...,\epsilon_{L}\}$ obtained by solving Eq.(6) when $\tau=4$.
Based on Figure 5, we observe that only a few layers are the most detrimental to safety, indicating that these few layers lack robustness to activation approximations. When $\tau=4$, the harmfulness of the model has reached its peak. As $\tau$ increases, the harmfulness hardly increases significantly, but the model utility decreases (PPL is increasing).
To show the noise distribution at this time, we present the layer-wise noise when $\tau=4$ in Figure 7. It can be seen that all the noise is distributed in the first $\tau$ layers, proving that these early layers are the least robust in safety. Therefore, we define the first $\tau$ layers as the sensitive layers.
@@
We can observe that before activation approximations, there exists a clear distinction between benign and harmful activations, with an average cosine similarity of $0.52$ among harmful activations. After activation approximations, harmful prompts become more scattered and diverse, overlapping significantly with benign regions, resulting in an average cosine similarity of $0.40$ among harmful activations. This shift, caused by activation approximations, enables harmful prompts to generate model internal states that resemble those of benign prompts, thereby bypassing safety alignment.